# LEXA: Legal Case Retrieval via Graph Contrastive Learning with Contextualised LLM Embeddings

### v2

@@ 2.2 Graph Contrastive Learning
Recently, contrastive learning is also applied in graph domain to fully leverage the graph structural information and explore the unlabelled graph data. The main difference between contrastive learning in the graph domain and other domains is the augmentation methods of graphs focusing on the aspects of graph structure such as node, edge or subgraph, where other domains have no graph structure. Specifically, there are various graph augmentation methods used for graph contrastive learning. Edge dropping is firstly proposed by Rong et al [DropEdge]to remove a certain portion of edges from the original graph. Similarly, node dropping [GRAND, GraphCL] refers to removing nodes with a certain probability. Besides of the graph structure augmentation such as removing nodes and edges, graph feature masking [BGRL, GraphCL] is also a popular graph augmentation method to perform efficient graph augmenting or corrupting by randomly mask the features of nodes or edges. Considering graph can be divided into subgraph, another graph augmentation approach called subgraph cropping is to crop out subgraph from the original graph then use the subgraph itself or the remaining graph as samples for contrastive learning (e.g.,  [GraphCL, GraphCrop]).
@@ 2.3 LLM Embeddings
Large language models (LLMs) have demonstrated outstanding performance in encoding text into high-quality embeddings for a wide range of natural language processing tasks, such as information retrieval and question answering. Their impressive encoding capability arises from the ability to model complex linguistic patterns and contextual dependencies through deep transformer architectures pretrained on large-scale corpora. Various LLM-based embedding models have been proposed to enhance semantic representation across diverse tasks. The recently proposed Multilingual Massive Text Embedding Benchmark (MMTEB) [mmteb] offers a comprehensive evaluation framework for assessing text embedding models across multiple tasks, such as classification, retrieval, and clustering. Legal retrieval is a domain-specific task included in MMTEB, where several state-of-the-art text embedding LLMs have achieved notable performance. voyage-law-2111https://blog.voyageai.com/2024/04/15/domain-specific-embeddings-and-retrieval-legal-edition-voyage-law-2/ ranks first in the MMTEB legal retrieval task by leveraging high-quality legal tokens, and a novel contrastive learning algorithm. E5-Mistral-7B-Instruct[e5mistral] is another top-performing model that utilizes proprietary LLMs to generate diverse synthetic data for training. OpenAI’s text-embedding-3-large [^2] https://platform.openai.com/docs/guides/embeddings is a high-quality general purpose embedding model, though it is primarily optimized for English and not specifically designed for multilingual retrieval tasks such as those in MMTEB. Additionally, Qwen3-Embedding-8B[qwen3embedding], trained using an innovative multi-stage pipeline that combines large-scale unsupervised pretraining with supervised fine-tuning on high quality datasets, demonstrates strong performance across both general purpose and domain-specific retrieval tasks. These advancements highlight the growing effectiveness and specialization of LLM-based embedding models in both general and domain-specific retrieval scenarios, particularly within the legal domain.
@@ 4 Method
Figure 1: The framework of LEXA. Given legal case $c$, the legal fact and the legal issue sections are converted into case graph. The case graph is processed by L layers of EUGAT and a Readout function to obtain an overall case graph representation. The whole framework is trained with the contrastive loss with positive, negative and augmented samples.
@@ 4.5 Graph Augmentation Contrastive Learning Objective Function
Specifically, in the LEXA model, graph augmentation is applied to both positive and negative cases to provide additional training signals, addressing the challenges posed by limited and costly legal data annotation. In Equation 18, the augmented graph of relevant cases $d^{+}$ is denoted as $d_{aug}^{+}$ while the augmented graph of easy negative sample $d^{easy-}$ is denoted as $d_{aug}^{easy-}$. During the graph contrastive learning process, only positive and easy negative samples are augmented; hard negative samples $d^{hard-}$ are used in their original form without augmentation. This design is motivated by experimental observations indicating that a limited number of hard negative samples is sufficient for effective training, whereas an excessive number can lead to notable performance degradation. Furthermore, for each case graph, only a single augmentation strategy is employed to generate an augmented graph. This decision is motivated by the empirical finding that using multiple augmentation methods simultaneously results in reduced performance. Additionally, during training, either augmented positive samples or augmented easy negative samples are used individually, as combining both was found to yield suboptimal results. The complete model pipeline is illustrated in Fig. 1.
@@ 5 Experiments
  • RQ4: How effective is prompt-based LLM features encoding?
  • RQ5: How does graph contrastive learning help with LCR in LEXA?
  • RQ6: How do hyper-parameter settings affect LEXA?
@@ 5.5 Effectiveness of Different LLM for Feature Encoding (RQ4)
In this section, the effectiveness of various LLMs for feature encoding is evaluated. Both E5-Mistral-7B-Instruct and Qwen3-Embedding-8B are employed to encode relation triplets into node and edge feature embeddings. Specifically, Qwen3-Embedding-8B is adopted as the feature encoder in LEXA.
@@ 5.7 Effectiveness of Graph Augmentation for Graph Contrastive Learning (RQ6)
5.8 Parameter Sensitivity (RQ7)

### v3

@@ 2.2 Graph Contrastive Learning
Recently, contrastive learning is also applied in graph domain to fully leverage the graph structural information and explore the unlabelled graph data. The main difference between contrastive learning in the graph domain and other domains is the augmentation methods of graphs focusing on the aspects of graph structure such as node, edge or subgraph, where other domains have no graph structure. Specifically, there are various graph augmentation methods used for graph contrastive learning. Edge dropping is firstly proposed by Rong et al [DropEdge]to remove a certain portion of edges from the original graph. Similarly, node dropping [GRAND, GraphCL] refers to removing nodes with a certain probability. Besides of the graph structure augmentation such as removing nodes and edges, graph feature masking [BGRL, GraphCL] is also a popular graph augmentation method to perform efficient graph augmenting or corrupting by randomly mask the features of nodes or edges. Considering graph can be divided into subgraph, another graph augmentation approach called subgraph cropping is to crop out subgraph from the original graph then use the subgraph itself or the remaining graph as samples for contrastive learning (e.g.,  [GraphCL, GraphCrop, puma, tnt-ood]).
@@ 2.3 LLM Embeddings
Large language models (LLMs) have demonstrated outstanding performance in encoding text into high-quality embeddings for a wide range of natural language processing tasks, such as information retrieval and question answering. Their impressive encoding capability arises from the ability to model complex linguistic patterns and contextual dependencies through deep transformer architectures pretrained on large-scale corpora. Various LLM-based embedding models have been proposed to enhance semantic representation across diverse tasks. The recently proposed Multilingual Massive Text Embedding Benchmark (MMTEB) [mmteb] offers a comprehensive evaluation framework for assessing text embedding models across multiple tasks, such as classification, retrieval, and clustering. Legal retrieval is a domain-specific task included in MMTEB, where several state-of-the-art text embedding LLMs have achieved notable performance. voyage-law-2111https://blog.voyageai.com/2024/04/15/domain-specific-embeddings-and-retrieval-legal-edition-voyage-law-2/ ranks first in the MMTEB legal retrieval task by leveraging high-quality legal tokens, and a novel contrastive learning algorithm. E5-Mistral-7B-Instruct[e5mistral] is another top-performing model that utilizes proprietary LLMs to generate diverse synthetic data for training. OpenAI’s text-embedding-3-large [^2] https://platform.openai.com/docs/guides/embeddings is a high-quality general purpose embedding model, though it is primarily optimized for English and not specifically designed for multilingual retrieval tasks such as those in MMTEB. Additionally, Qwen3-Embedding-8B[qwen3embedding], trained using an innovative multi-stage pipeline that combines large-scale unsupervised pretraining with supervised fine-tuning on high quality datasets, demonstrates strong performance across both general purpose and domain-specific retrieval tasks. Specifically, some LLM-embedding–based legal retrieval methods, such as UQLegalAI@COLIEE2025 [uqlegalai], and legally fine-tuned large language models, such as ReaKase-8B [reakase], have been proposed to enhance semantic understanding of legal texts. These advancements highlight the growing effectiveness and specialization of LLM-based embedding models in both general and domain-specific retrieval scenarios, particularly within the legal domain.
@@ 4 Method
Figure 1: Given a legal case $c$, the legal fact and legal issue sections are extracted and transformed into structured case graphs using LEXA-8B and case graph generation (TACG). Each case graph is encoded by L stacked Edge-Updated Graph Attention Network (EUGAT) layers to model semantic and relational dependencies. A readout function aggregates node representations into graph-level embeddings, which are combined to form the final case representation $h_{c}$. The model is trained end-to-end with a contrastive learning objective incorporating positive, easy and hard negative, and augmented samples to improve discriminative power and robustness.
@@ 4.5 Graph Augmentation Contrastive Learning Objective Function
Specifically, in the LEXA model, graph augmentation is applied to both positive and negative cases to provide additional training signals, addressing the challenges posed by limited and costly legal data annotation. In Equation 18, the augmented graph of relevant cases $d^{+}$ is denoted as $d_{aug}^{+}$ while the augmented graph of easy negative sample $d^{easy-}$ is denoted as $d_{aug}^{easy-}$. For each case graph, only a single augmentation strategy is employed to generate an augmented graph. This decision is motivated by the empirical finding that using multiple augmentation methods simultaneously results in reduced performance. Additionally, during training, either augmented positive samples or augmented easy negative samples are used individually, as combining both was found to yield suboptimal results.
@@ 5 Experiments
  • RQ4: How different LLMs affect the performance
  • RQ5: How effective is prompt-based LLM features encoding?
  • RQ6: How does graph contrastive learning help with LCR in LEXA?
  • RQ7: How does graph sparsity affect retrieval performance in LEXA?
  • RQ8: How do hyper-parameter settings affect LEXA?
@@ 5.5 Effectiveness of Different LLM for Feature Encoding (RQ4)
In this section, the effectiveness of various LLMs for feature encoding is evaluated on both datasets across all evaluation metrics, with results reported in Table 7. Both E5-Mistral-7B-Instruct and Qwen3-Embedding-8B are employed to encode relation triplets into node and edge feature embeddings. Specifically, Qwen3-Embedding-8B is adopted as the feature encoder in LEXA.
@@ 5.7 Effectiveness of Graph Augmentation for Graph Contrastive Learning (RQ6)
5.8 Impact of Graph Sparsification via Edge Pruning (RQ7)
Table 11: Impact of graph sparsification via edge pruning.
Methods COLIEE2022
P@5 R@5 Mi-F1 Ma-F1 MRR@5 MAP NDCG@5
80% Edge 39.6 $\pm$0.4 47.1 $\pm$0.5 43.0 $\pm$0.5 47.7 $\pm$0.5 71.0 $\pm$0.3 68.4 $\pm$0.3 74.5 $\pm$0.1
90% Edge 40.5 $\pm$0.5 48.1 $\pm$0.6 43.9 $\pm$0.6 48.6 $\pm$0.7 72.6 $\pm$2.2 70.0 $\pm$1.6 75.7 $\pm$1.5
100% Edge 43.7 $\pm$0.2 51.9 $\pm$0.2 47.5 $\pm$0.2 52.4 $\pm$0.3 77.0 $\pm$0.5 74.7 $\pm$0.9 79.3 $\pm$0.8
Methods COLIEE2023
P@5 R@5 Mi-F1 Ma-F1 MRR@5 MAP NDCG@5
80% Edge 22.2 $\pm$0.1 41.3 $\pm$0.1 28.9 $\pm$0.1 29.9 $\pm$0.1 47.9 $\pm$0.3 46.1 $\pm$0.3 51.7 $\pm$0.3
90% Edge 22.4 $\pm$0.4 41.7 $\pm$0.2 29.2 $\pm$0.2 30.3 $\pm$0.2 48.6 $\pm$0.4 46.8 $\pm$0.6 51.4 $\pm$0.5
100% Edge 22.6 $\pm$0.4 41.9 $\pm$0.8 29.3 $\pm$0.6 30.5 $\pm$0.7 48.7 $\pm$0.7 47.0 $\pm$0.7 52.6 $\pm$0.8
To evaluate how reducing graph density affects retrieval effectiveness, three experiments with different edge pruning ratios are conducted on both datasets, as shown in Table 11. Specifically, the 80% edge setting retains 80% of the edges from the original TACG graph, the 90% edge setting retains 90% of the edges, and the 100% edge setting corresponds to the original, unpruned graph. Overall, the results indicate that reducing graph density through edge pruning leads to a steady degradation in retrieval effectiveness across all evaluation metrics. The unpruned graph consistently achieves the best performance on both datasets, outperforming pruned variants in terms of P@5, F1, MAP, MRR@5, and NDCG@5. Moderate edge pruning (90%) results in noticeable but limited performance drops, while more aggressive pruning (80%) causes further degradation. This trend suggests that excessive edge removal disrupts important structural and semantic relationships that are essential for accurate legal case retrieval. Although the performance gap between pruned and unpruned graphs is smaller on COLIEE2023 than on COLIEE2022, the fully connected graph remains consistently superior. These findings demonstrate that preserving rich graph connectivity is crucial for stable ranking quality across different dataset scales, and naive edge pruning introduces a clear accuracy–efficiency trade-off.
5.9 Parameter Sensitivity (RQ8)
@@ 6 Conclusion
Declarations
Funding declaration This work is supported by Australian Research Council CE200100025, DP230101196, DP230101753 and DE250100919.
Ethics declaration Not applicable.